OpenStudy (telena):

A man drove a car 191.25 miles. The cars mileage for the drive was 25.5 miles per gallon. How many gallons of gas did the car use?

OpenStudy (maddielb):

Well if it went 191.25 miles and it gets 25.5 miles per gallon, you just divide 191.25 by 25.5 and get 7.5 gallons

OpenStudy (telena):

That isnt the right answer though 🙃

OpenStudy (telena):

I tried it marked it incorrect

OpenStudy (maddielb):

Uhh...

OpenStudy (maddielb):

Did you put just 7.5 or 7.5 gallons

OpenStudy (telena):

No, theres options in all fractions and i put 7 1/5

OpenStudy (maddielb):

As a fraction it would be 75/10 I believe

OpenStudy (danjs):

You are converting miles , into the gallons used from the given ratio. \[\frac{ 191.25~mi }{ 1 }*\frac{ 1~gal }{ 25.5~mi }=7.5~~gal\]

OpenStudy (danjs):

you want a fraction.. \[\frac{ 7.5 }{ 1 }*\frac{ 10 }{ 10 }=\frac{ 75 }{ 10 } = \frac{ 15 }{ 2 }\]

OpenStudy (telena):

Theres the options

OpenStudy (danjs):

7.5 is the same as seven and a half

OpenStudy (telena):

Alright, thanks

OpenStudy (masonsurfs):

Hi, Im pretty sure 7.5 is correct